OpenStudy (anonymous):

use the given conditions to write an equation for the line in point slope form, slope-intercept form, and standard form passing through (-4,6) and parallel to the line whose equation is Y= -2/3x +1/3 (HELP PLEASE)

OpenStudy (anonymous):

what is said about the gradients of parallel slopes?

OpenStudy (anonymous):

@mickey4691 IT SAYS NOTHING ABOUT THE GRADIENTS

OpenStudy (anonymous):

if two lines are parallel, that means they have the same gradient

OpenStudy (anonymous):

use the given conditions to write an equation for the line in point slope form, slope-intercept form, and standard form passing through (-4,6) and parallel to the line whose equation is Y= -2/3x +1/3 (HELP PLEASE)

OpenStudy (anonymous):

try to understand what's going on first

OpenStudy (anonymous):

that's all it says @mick

OpenStudy (anonymous):

the point (-4, 6) has the same gradient as the other equation because the lines are parallel

OpenStudy (anonymous):

oh ok I see what you're saying @mickey4691 sorry about that this isn't my best subject

OpenStudy (anonymous):

that means my y=6, x=-4 and m= -2/3 so i have to find c to make y=mx+c

OpenStudy (anonymous):

therefore i have 6= (-2/3)*(-4) + c find for c and rewrite the equation as a slope form

OpenStudy (anonymous):

so is that point slope form

OpenStudy (anonymous):

??? @mickey4691

OpenStudy (anonymous):

i have to find for c. c becomes 10/3 therefore the point slope form becomes y=-2/3x + 10/3

OpenStudy (anonymous):

got it so what's the slope-intercept form? @mickey4691